a line has equation y=3x+7 and passes through the point(h,h+15).calculate the value of h?
Found 2 solutions by sachi, Cromlix:
Answer by sachi(548)

y=3x+7 and passes through the point(h,h+15)
so h+5=3h+7
or 3h-h=5-7
or 2h=-2
or h=-1

Answer by Cromlix(4381)

y = 3x + 7
Substituting h for x
Substituting h + 15 for y
h + 15 = 3(h) + 7
h + 15 = 3h + 7
-2h = -8
h = 4
x = 4
y = 19
Hope this helps.
:-)